Solve.
ln(3x+8)-6=3

Answer by lwsshak3(11628)

Solve.
ln(3x+8)-6=3
**
ln(3x+8)-6=3
ln(3x+8)=3+6=9
convert to exponential form:base(e) raised to log of number(9)=number(3x+8)
e^9=3x+8
3x=e^9-8=8103.0839-8=8095.0839
x=8095.0839/3
x=2698.3613
Check:
ln(3x+8)
ln(3*2698.3613+8)
ln(8103.0839)=9